OpenStudy (nikolas):

Prove that (1-tanx)/(1-cotx) = -tanx I've got the first step, that (1-sinx/cosx)/(1-cosx/sinx) = -tanx, but I'm not sure what to do next. I've been told that the next step is "(cosx-sinx)/cosx * (sinx)/ (sinx - cosx) " but I'm not sure what was done to get there.

OpenStudy (fwizbang):

If it were me, I'd start differently. Pull a factor of tan x out pf the numerator to get tan x * (1/tan x -1)/(1 - cot x) and then use cot x = 1/ tan x.

OpenStudy (nikolas):

Sorry, I'm a little confused about what you did there - how does (1-tanx) equal tanx*(1/tanx-1)?

OpenStudy (fwizbang):

\[1 - \tan x = (1- \tan x ) * {\tan x \over \tan x } = ({1 \over \tan x} -1)* \tan x\]

OpenStudy (turingtest):

though fwzbangs way is faster anyway, here's what you wanted:\[{1-\tan x\over1-\cot x}={\cos x\sin x\over\cos x\sin x}\cdot{{1-\frac{\sin x}{\cos x}}\over1-\frac{\cos x}{\sin x}}={\sin x\cos x-\sin^2x\over\sin x\cos x-\cos^2x}\]\[{\sin x\over\cos x}\cdot{\cos x-\sin x\over\sin x-\cos x}\]

OpenStudy (turingtest):

so the trick was to multiply by\[\sin x\cos x\over\sin x\cos x\]

OpenStudy (nikolas):

Thank you both, I've got it now. I'll give best response to TuringTest for answering the original question but I really appreciate the explanation, fwizbang!